# Asymmetric Graph Error Control with Low Complexity in Causal Bandits

### v1

@@ VI-B Performance for Stationary Causal Bandits
Next, we consider the proposed CSL-UCB scheme with optimized intervention design in causal bandits. Since the goal is to maximize the cumulative reward, we start by revealing the impact of FN errors on reward estimation. In Fig. 4, the normalized empirical reward error is plotted as a function of time for two cases: with or without FN errors. Without FN errors, the reward estimation error decreases monotonically as more samples are collected, which matches the analysis in Section III-C. With the presence of FN errors, the estimation error does not decrease as more samples are collected. On average, the estimation error is $71.9\%$ lower in the absence of FN errors. The result is consistent with our claim that FN errors have a significant impact on reward maximization in causal bandits. The result also validates our choice to minimize the FN rate in causal graph identification.
@@
To evaluate the CSL-UCB scheme, we focus on graphs with size $N=10$ and simulate causal bandits over a horizon of $T=1500$ time steps. For comparison, we first consider the vanilla UCB algorithm [22], which selects an intervention in each step as
@@
where $\alpha^{\prime}$ is the parameter that controls the exploration level. Notice that the vanilla UCB algorithm does not exploit the causal structure and its sample complexity scales exponentially as $2^{N}$. Another comparison scheme we consider is GOLEM [7] based decision-making for MABs, denoted as the GOLEM-MAB scheme. The horizon is divided into two parts, such that causal structure identification is conducted in the first part, while rewards are collected in the second part by exploiting the learned causal graph. Note that since GOLEM identifies the whole graph, intervention is required to be the all-zeros vector for learning $\bm{B}$ and the all-ones for learning $\bm{B}^{\prime}$. Lastly, we employ the LinSEM-TS scheme proposed in [17], which provides an upper bound on performance, as it utilizes knowledge of the causal topology.
Figure 5 plots the cumulative regret as a function of time for four different algorithms and the associated probability of selecting the optimal intervention is provided in Figure 6. The vanilla UCB algorithm gains information about the optimal intervention after exploring every possible intervention at $t=2^{N}=1024$. By the end of the horizon, vanilla UCB selects the optimal intervention $48.9\%$ of the time. GOLEM-MAB tries to identify the causal graph in the first $600$ steps, and exploits this knowledge to achieve an optimal intervention selection ratio of $71.0\%$. The value of $t=600$ was selected via trial and error to optimize the performance of GOLEM-MAB. The proposed CSL-UCB scheme gains and leverages causal knowledge in an alternating manner, yielding an optimal intervention selection ratio of $78.2\%$. With topology knowledge, LinSEM-TS is able to select the optimal intervention with probability $94.9\%$. In terms of cumulative regret, the CSL-UCB algorithm achieves a $90.8\%$ improvement compared with vanilla UCB and a $85.8\%$ improvement compared with GOLEM-MAB. The regret gap between LinSEM-TS and CSL-UCB is relatively small.
Figure 5: Cumulative regret as a function of time steps, stationary bandits.
Figure 6: Optimal percentage as a function of time steps, stationary bandits.
An interesting observation is that, after $600$ time steps for graph identification ($300$ for each, $\bm{B}$ and $\bm{B}^{\prime}$), the graphs estimated by GOLEM are close to the ground truth with respect to the SHD. However, even rejecting one existing edge (one false negative error) could result in a significantly different expected reward. The relatively poor performance of GOLEM-MAB is not due to its generic graph identification ability, but the equal treatment of all errors and its low sample efficiency. We emphasize again that causal bandit optimization and causal graph identification are two different problems.
@@ VI-C Performance for Non-stationary Causal Bandits
Lastly, we evaluate the reward-earning ability of the proposed CSL-UCB-CD scheme in non-stationary causal bandits. Structural changes occur twice at $t=1000,2000$, where a random subset of the causal mechanisms are regenerated. For comparison, the vanilla UCB algorithm can adjust its estimate of the rewards after structural changes using newly collected samples. Knowledge of the current causal structure is given to LinSEM-TS resulting in an upper bound on performance. The GOLEM-MAB algorithm combined with change detection is denoted as GOLEM-MAB-CD, which must re-learns the whole causal graph when a change is detected.
In Fig. 7, the cumulative regret is plotted as a function of time. The percentage of selecting the optimal intervention as a function of time is provided in Fig. 8. The cumulative regret achieved by the CSL-UCB-CD scheme is $83.9\%$ lower than the vanilla UCB scheme and $83.1\%$ lower than GOLEM-MAB-CD. As an upper bound on performance, LinSEM-TS achieves the lowest cumulative regret, while the gap between LinSEM-TS and CSL-UCB-CD is again relatively small.
Figure 7: Cumulative regret as a function of time steps, non-stationary bandits.
Figure 8: Optimal percentage as a function of time steps, non-stationary bandits.
The vanilla UCB scheme performs the worst, not due to structure changes but rather, due to the lack of samples. Averaged over time, the optimal intervention ratio of the vanilla UCB scheme is only $8.7\%$. As illustrated in Fig. 8, the optimal percentage of GOLEM-MAB-CD drops to $0\%$ after each change-point, which indicates that it is able to detect the changes and start to re-learn the causal graph. However, as learning the whole graph requires a large amount of samples, GOLEM-MAB-CD is only able to select the optimal intervention $17.5\%$ of the time. With the latest structure knowledge, LinSEM-TS does not suffer from structure changes, and selects the optimal intervention $89.8\%$ of the time. The CSL-UCB-CD scheme combines sub-graph change detection with sub-graph learning, and achieves the optimal intervention ratio of $67.4\%$.
@@ VII Conclusions
In this paper, we investigate the causal bandit problem with no knowledge of the causal graph topology and the interventional distribution. Distinguishing between the two types of error, a sub-graph learning-based causal identification approach is proposed, which makes efficient use of the limited data. Due to learning sub-graphs versus entire graphs, the computational complexity of the proposed method is significantly lower than traditional methods. Analysis reveals that optimal intervention design is much more sensitive to false negative errors in graph identification. This feature is directly taken into consideration for the sub-graph learning algorithm. To balance exploration and exploitation, we analyze and propose an uncertainty bound. Further, sub-graph level change detection mechanism is proposed, which enables the proposed approach to adapt to non-stationary bandits. Numerical results show that the overall approach is able to identify the optimal intervention with much lower sample and computational complexity than previous methods. Furthermore, performance is improved over prior approaches due to the direct consideration of false negative errors. The proposed approach requires $67\%$ fewer samples to learn the causal structure and achieves a lower cumulative regret by $85\%$, compared to the existing schemes.
@@ Appendix A Proof of Proposition 2
$Z_{1}$ $=\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\epsilon_{i},$
$Z_{2}$ $=\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\hat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k},$
@@ Appendix B Proof of Proposition 3
$\bm{X}^{1:t}_{i}(a_{i})$ $=\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}^{T}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})+\bm{0}^{T}\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i})$ (71)
$=\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}}^{T}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i})$ (72)
@@
$\mathbb{E}\bigl{[}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\mathcal{S}}\big{|}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})\bigr{]}=\mathbb{E}\bigl{[}\bigl{[}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})\bm{X}^{1:t}_{\mathcal{S}}(a_{i})^{T}\bigr{]}^{-1}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})$
$\hskip 38.0pt\cdot(\bm{X}^{1:t}_{\mathcal{S}}(a_{i})^{T}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}}+\bm{\epsilon}^{1:t}_{i}(a_{i})^{T})\big{|}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})\bigr{]}$ (73)
@@
With the substitution $\mathcal{Q}=\mathcal{P}_{i}(a_{i})$, $\mathcal{S}=\hat{\mathcal{P}}_{i}(a_{i})$, we obtain the effect of including non-existent parents (false positive error). Equation (74) holds because the exogenous variable $\epsilon_{i}$ is independent of the variables associated with the nodes in $\mathcal{S}=\hat{\mathcal{P}}_{i}(a_{i})$. Without a false negative error, $\hat{\mathcal{P}}_{i}(a_{i})$ should not contain node $i$ and its descendants.
@@
$\bm{X}^{1:t}_{i}(a_{i})=\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}^{T}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})+\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}\backslash\mathcal{Q}}^{T}\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i}).$ (75)
@@
$\mathbb{E}\bigl{[}\bigl{[}\hat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\mathcal{Q}}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}=\mathbb{E}\bigl{[}\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})^{T}\bigr{]}^{-1}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})$
$\hskip 24.0pt\bigl{(}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})^{T}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}+\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})^{T}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}\backslash\mathcal{Q}}+$
$\hskip 138.0pt\bm{\epsilon}^{1:t}_{i}(a_{i})^{T}\bigr{)}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}$ (76)
$\quad=\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}+\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})^{T}\bigr{]}^{-1}\mathbb{E}\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\cdot$
$\hskip 84.0pt\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})^{T}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}\backslash\mathcal{Q}}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}.$ (77)
Note that having FN errors is equivalent to have $\hat{\mathcal{P}}_{i}(a_{i})\subset\mathcal{P}_{i}(a_{i})$. Thus with the substitution $\mathcal{Q}=\hat{\mathcal{P}}_{i}(a_{i})$, $\mathcal{S}=\mathcal{P}_{i}(a_{i})$, we obtain the effect of FN errors (rejecting actual edges). ∎

### v2

@@ VI-B Performance for Stationary Causal Bandits
Next, we consider the proposed CSL-UCB scheme with optimized intervention design in causal bandits. Since the goal is to maximize the cumulative reward, we start by revealing the impact of FN errors on reward estimation. To measure the impact, the estimation errors are classified into two groups based on whether FN error exists in the estimated graph. In Fig. 6, the normalized empirical reward error is plotted as a function of time for two cases: with or without FN errors. Without FN errors, the reward estimation error decreases monotonically as more samples are collected, which matches the analysis in Section III-C. With the presence of FN errors, the estimation error does not decrease as more samples are collected. On average, the estimation error is $62.8\%$ lower in the absence of FN errors. The result is consistent with our claim that FN errors have a significant impact on reward maximization in causal bandits. The result also validates our choice to minimize the FN rate in causal graph identification.
@@
To evaluate the CSL-UCB scheme, we focus on graphs with size $N=10$ and simulate causal bandits over a horizon of $T=1500$ time steps. To alleviate computation cost, the graphs are updated with new collected samples every $20$ time slots. For comparison, we first consider the vanilla UCB algorithm [29], which selects an intervention in each step as
@@
where $\alpha^{\prime}$ is the parameter that controls the exploration level. Notice that the vanilla UCB algorithm does not exploit the causal structure and its sample complexity scales exponentially as $2^{N}$ [46]. Another scheme we consider for comparison is the GOLEM [10] algorithm based decision-making for MABs, denoted as the GOLEM-MAB scheme. The horizon is divided into two parts, such that causal structure identification is conducted in the first part, while rewards are collected in the second part by exploiting the learned causal graph. Note that since GOLEM identifies the whole graph, intervention should be fixed as $\bm{a}=\bm{0}$ for learning $\bm{B}$ and $\bm{a}=\bm{1}$ for learning $\bm{B}^{\prime}$. Lastly, we employ the LinSEM-TS scheme proposed in [20], which provides an upper bound on performance, as it utilizes knowledge of the causal topology.
Figure 7 plots the cumulative regret as a function of time for four different algorithms and the associated probability of selecting the optimal intervention is provided in Figure 8. The vanilla UCB algorithm gains information about the optimal intervention after exploring every possible intervention at $t=2^{N}=1024$. By the end of the horizon, vanilla UCB selects the optimal intervention $47.2\%$ of the time. GOLEM-MAB tries to identify the causal graph in the first $600$ steps, and exploits this knowledge to achieve an optimal intervention selection ratio of $71.0\%$. The value of $t=600$ was selected via trial and error to optimize the performance of GOLEM-MAB. The proposed CSL-UCB scheme gains and leverages causal knowledge in an alternating manner, yielding an optimal intervention selection ratio of $79.0\%$. With topology knowledge, LinSEM-TS is able to select the optimal intervention with probability $94.2\%$. In terms of cumulative regret, the CSL-UCB scheme achieves a $91.6\%$ improvement compared with vanilla UCB and an $86.8\%$ improvement compared with GOLEM-MAB. The LinSEM-TS algorithm can further reduce $86.6\%$ of the cumulative regret the CSL-UCB scheme achieves when the causal structure is fully known.
Figure 7: Cumulative regret of stationary bandits as a function of time steps.
Figure 8: Optimal percentage of stationary bandits as a function of time steps.
An interesting observation is that, after $600$ time steps for graph identification ($300$ for each, $\bm{B}$ and $\bm{B}^{\prime}$), the graphs estimated by GOLEM are close to the ground truth with respect to the SHD. However, even rejecting one existing edge (one false negative error) could result in a significantly different expected reward. Since edges are highly possible to be part of the causal flows that influence the reward node, mistakenly rejecting an edge can cut off one or multiple relevant causal flows and thus significantly impact reward estimation. The relatively poor performance of GOLEM-MAB is not due to its generic graph identification ability, but the equal treatment of all errors and its low sample efficiency. We emphasize again that causal bandit optimization and causal graph identification are two different problems.
@@ VI-C Performance for Non-stationary Causal Bandits
Lastly, we evaluate the reward-earning ability of the proposed CSL-UCB-CD scheme in non-stationary causal bandits. Structure change occurs twice at $t\in\{1000,2000\}$, where a random subset of the causal mechanisms are regenerated. For comparison, the vanilla UCB algorithm can adjust its estimate of the rewards after structure changes using newly collected samples. The GOLEM-MAB algorithm combined with a change detection mechanism is denoted as GOLEM-MAB-CD, which must re-learns the whole causal graph when a change is detected. The LinSEM-TS algorithm initially has full knowledge of the causal structure, but is unaware of the upcoming changes.
In Fig. 9, the cumulative regret is plotted as a function of time. The percentage of selecting the optimal intervention as a function of time is provided in Fig. 10. On average, the change detection mechanism is able to detect $89.3\%$ of the sub-graph changes, with an average delay of $13.1$ slots. The cumulative regret achieved by the CSL-UCB-CD scheme is $83.0\%$ lower than vanilla UCB, $83.2\%$ lower than GOLEM-MAB-CD and $73.5\%$ lower than LinSEM-TS. Although LinSEM-TS obtains the lowest regret before structure change, its performance degrades dramatically when the causal knowledge becomes incorrect.
Figure 9: Cumulative regret of non-stationary bandits as a function of time steps. Structure change occurs twice, as marked by the dashed lines.
Figure 10: Optimal percentage of non-stationary bandits as a function of time steps. Structure change occurs twice, as marked by the dashed lines.
The vanilla UCB scheme performs the worst, not due to structure changes but rather, due to the lack of samples. Averaged over time, the optimal intervention ratio of the vanilla UCB scheme is only $7.0\%$. As illustrated in Fig. 10, the optimal percentage of GOLEM-MAB-CD drops to $0\%$ after each change-point, which indicates that it is able to detect the changes and start to re-learn the causal graph. However, as learning the whole graph requires a large amount of samples, GOLEM-MAB-CD is only able to select the optimal intervention $18.3\%$ of the time. Since LinSEM-TS suffers from incorrect causal knowledge caused by structure changes, it selects the optimal intervention only $32.6\%$ of the time. The CSL-UCB-CD scheme combines sub-graph change detection with sub-graph learning, and select the optimal intervention $63.8\%$ of the time.
@@ VII Conclusions
In this paper, we investigate the causal bandit problem with no knowledge of the causal graph topology and the interventional distribution. Distinguishing between the two types of error, a sub-graph learning-based causal identification approach is proposed, which makes efficient use of the limited data. Due to learning sub-graphs versus entire graphs, the complexity of the proposed method is significantly lower than traditional methods. Analysis reveals that optimal intervention design is much more sensitive to false negative errors in graph identification. This feature is directly taken into consideration for the sub-graph learning algorithm. To balance exploration and exploitation, we analyze and propose an uncertainty bound. Further, sub-graph level change detection mechanism is proposed, which enables the proposed approach to adapt to non-stationary bandits. Numerical results show that the overall approach is able to identify the optimal intervention with much lower sample and computational complexity than previous methods. Furthermore, performance is improved over prior approaches due to the direct consideration of false negative errors. The proposed approach requires $67\%$ fewer samples to learn the causal structure and achieves a lower cumulative regret by $85\%$, compared to the existing schemes.
@@ Appendix A Proof of Proposition 2
$Z_{1}$ $=\sum_{l\not\in\mathcal{A}_{i}(a_{i})}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{li}\epsilon_{l}-\tilde{\epsilon}_{i},$
$Z_{2}$ $=\sum_{k\in\mathcal{A}_{i}(a_{i})}\bigl{(}\bigl{[}\widehat{\bm{C}}_{\bm{a}}^{t}\bigr{]}_{ki}-\left[\bm{C}_{\bm{a}}\right]_{ki}\bigr{)}\epsilon_{k},$
@@ Appendix B Proof of Proposition 3
$\bm{X}^{1:t}_{i}(a_{i})$ $={\left[\bm{B}_{\bm{a}}\right]}^{\top}_{i,\mathcal{Q}}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})+{\bm{0}}^{\top}\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i})$ (73)
$={\left[\bm{B}_{\bm{a}}\right]}^{\top}_{i,\mathcal{S}}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i})$ (74)
@@
$\mathbb{E}\bigl{[}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\mathcal{S}}\big{|}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})\bigr{]}=\mathbb{E}\bigl{[}\bigl{[}\bm{X}^{1:t}_{\mathcal{S}}(a_{i}){\bm{X}^{1:t}_{\mathcal{S}}(a_{i})}^{\top}\bigr{]}^{-1}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})$
$\hskip 38.0pt\cdot({\bm{X}^{1:t}_{\mathcal{S}}(a_{i})}^{\top}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}}+{\tilde{\bm{\epsilon}}^{1:t}_{i}(a_{i})}^{\top})\big{|}\bm{X}^{1:t}_{\mathcal{S}}(a_{i})\bigr{]}$ (75)
@@
With the substitution $\mathcal{Q}=\mathcal{P}_{i}(a_{i})$, $\mathcal{S}=\widehat{\mathcal{P}}_{i}(a_{i})$, we obtain the effect of including non-existent parents (false positive error). Equation (76) holds because the exogenous variable $\tilde{\epsilon}_{i}$ is independent of the variables associated with the nodes in $\mathcal{S}=\widehat{\mathcal{P}}_{i}(a_{i})$. Without a false negative error, $\widehat{\mathcal{P}}_{i}(a_{i})$ should not contain node $i$ and its descendants.
@@
$\bm{X}^{1:t}_{i}(a_{i})={\left[\bm{B}_{\bm{a}}\right]}^{\top}_{i,\mathcal{Q}}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})+{\left[\bm{B}_{\bm{a}}\right]}^{\top}_{i,\mathcal{S}\backslash\mathcal{Q}}\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})+\bm{\epsilon}^{1:t}_{i}(a_{i}).$ (77)
@@
$\mathbb{E}\bigl{[}\bigl{[}\widehat{\bm{B}}_{\bm{a}}^{t}\bigr{]}_{i,\mathcal{Q}}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}=\mathbb{E}\bigl{[}\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i}){\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})}^{\top}\bigr{]}^{-1}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})$
$\hskip 24.0pt\bigl{(}{\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})}^{\top}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}+{\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})}^{\top}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}\backslash\mathcal{Q}}+$
$\hskip 138.0pt{\tilde{\bm{\epsilon}}^{1:t}_{i}(a_{i})}^{\top}\bigr{)}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}$ (78)
$\quad=\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{Q}}+\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i}){\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})}^{\top}\bigr{]}^{-1}\mathbb{E}\bigl{[}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\cdot$
$\hskip 84.0pt{\bm{X}^{1:t}_{\mathcal{S}\backslash\mathcal{Q}}(a_{i})}^{\top}\left[\bm{B}_{\bm{a}}\right]_{i,\mathcal{S}\backslash\mathcal{Q}}\big{|}\bm{X}^{1:t}_{\mathcal{Q}}(a_{i})\bigr{]}.$ (79)
Note that having FN errors is equivalent to have $\widehat{\mathcal{P}}_{i}(a_{i})\subset\mathcal{P}_{i}(a_{i})$. Thus with the substitution $\mathcal{Q}=\widehat{\mathcal{P}}_{i}(a_{i})$, $\mathcal{S}=\mathcal{P}_{i}(a_{i})$, we obtain the effect of FN errors (rejecting actual edges). ∎